OpenStudy (anonymous):

A charged particle is projected in the air such that its horizontal (x) and vertical (y) shifts are given by x(t) = 10t − t2 and y(t) = 6t, where x and y are in meters and time t is in seconds. What is the ratio of y to x at t = 4 seconds?Select one of the options below as your answer: A. 0.25 B. 2 C. 0.75 D. 0.85 E. 1

OpenStudy (whpalmer4):

You're simply finding the ratio of the two expressions when t = 4. \[\frac{y(t)}{x(t)} = \frac{6t}{10t-t^2} = \]